Consider a rectangle with perimeter 28 (units). Let the width of the rectangle be w (units) and let the Area of the region enclosed by the rectangle be A ( square units). Express A as a function of w and state the domain and range of the function.


1
Expert's answer
2021-10-26T17:44:57-0400

the width of the rectangle:

"l=(28-2w)\/2"


"A=wl=w(28-2w)\/2=14w-w^2"


domain: "w\isin (0,14)"


"A'(w)=14-2w=0"

"A_{max}=A(7)=14\cdot7- 7^2=49"

range: "A\isin (0,49]"